1.1 find parametric equations of the line that passes through the point P =(2, 0,-1) and is parallel to the vector n =<2, 1, 3>

1.2 find paramedic equations of the line that passes through the points A= (1, 2, - 3) and B =(7, 2, - 4).

1.3 find paramedic equations for the line of intersection of the planes - 5x + y - 2z =3 and 2x - 3y + 5z =-7


1
Expert's answer
2021-07-14T14:19:30-0400

QUESTION 1


Let "p.M\left(x,y,z\right)" another point on the straight line, the equation of which we are looking for.

Then, "\overrightarrow{PM}||\overrightarrow{n}, \,\,\,\overrightarrow{n}=\left(2,1,3\right)-\text{specified by condition}" .



"\left\{\begin{array}{l}
p.P\left(2,0,-1\right)\\[0.3cm]
p.M\left(x,y,z\right)
\end{array}\right.\rightarrow\overrightarrow{PM}\left(x-2,y,z+1\right)\\[0.3cm]
\overrightarrow{PM}||\overrightarrow{n}\rightarrow\frac{x-2}{2}=\frac{y}{1}=\frac{z+1}{3}=t\\[0.3cm]
\left\{\begin{array}{l}
x-2=2t\\[0.3cm]
y=t\\[0.3cm]
z+1=3t
\end{array}\right.\longrightarrow
\left\{\begin{array}{l}
x=2+2t\\[0.3cm]
y=t\\[0.3cm]
z=-1+3t
\end{array}\right."

ANSWER



"\left\{\begin{array}{l}
x=2+2t\\[0.3cm]
y=t\\[0.3cm]
z=-1+3t
\end{array}\right."

QUESRION 2


As we know, the parametric equation of the straight line has the form



"\left\{\begin{array}{l}
x=x_0+at\\[0.3cm]
y=y_0+bt\\[0.3cm]
z=z_0+ct
\end{array}\right.,\,\,\,\text{where}
\left\{\begin{array}{l}
p.M\left(x_0,y_0,z_0\right)-\text{any point on the line}\\[0.3cm]
\overrightarrow{n}=\left(a,b,c\right)-\text{direction vector}
\end{array}\right."

In our case, since by the problem statement we know two points "p.A=\left(1, 2, - 3\right)" and "p.B=\left(7, 2, - 4\right)" through which the straight line passes, then



"\overrightarrow{n}=\overrightarrow{AB}=\left(7-1,2-2,-4-\left(-3\right)\right)\\[0.3cm]
\boxed{\overrightarrow{n}=\left(6,0,-1\right)}"

As an "unknown" "p.M\left(x_0,y_0,z_0\right)" , I propose to take "p.A=\left(1,2,-3\right)" from the condition.

Then,



"\left\{\begin{array}{l}
x=1+6t\\[0.3cm]
y=2\\[0.3cm]
z=-3-t
\end{array}\right."

ANSWER



"\left\{\begin{array}{l}
x=1+6t\\[0.3cm]
y=2\\[0.3cm]
z=-3-t
\end{array}\right."

QUESTION 3


From the beginning, we find any point "p.M\left(x_0,y_0,z_0\right)" on the desired straight line, for this we solve the system of equations



"\left\{\begin{array}{l}
- 5x + y - 2z =3\to y=3+5x+2z\\[0.3cm]
2x - 3y + 5z =-7
\end{array}\right.\\[0.3cm]
\left\{\begin{array}{l}
y=3+5x+2z\\[0.3cm]
2x - 3\cdot\left(3+5x+2z\right) + 5z =-7
\end{array}\right.\\[0.3cm]
\left\{\begin{array}{l}
y=3+5x+2z\\[0.3cm]
2x - 9-15x-6z + 5z =-7
\end{array}\right.\\[0.3cm]
\left\{\begin{array}{l}
y=3+5x+2z\\[0.3cm]
-13x-z =2\to z=-2-13x
\end{array}\right.\\[0.3cm]
\left\{\begin{array}{l}
y=3+5x+2\cdot\left(-2-13x\right)\\[0.3cm]
z=-2-13x
\end{array}\right.\\[0.3cm]
\left\{\begin{array}{l}
y=-1-21x\\[0.3cm]
z=-2-13x\\[0.3cm]
x=Const
\end{array}\right.\\[0.3cm]"

Let "x=0\longrightarrow \boxed{p.M\left(0,-1,-2\right)}" .

As we know, from the canonical equation of the plane we can find the coordinates of the normal vector as follows



"l : ax+by+cz+d=0\to\overrightarrow{n_l}=\left(a,b,c\right)"

( more information : https://en.wikipedia.org/wiki/Normal_(geometry) )

In our case,



"\left\{\begin{array}{l}
l_1 :-5x + y - 2z =3\to\overrightarrow{n_{l_1}}=\left(-5,1,-2\right)\\[0.3cm]
l_2 :2x - 3y + 5z =-7\to\overrightarrow{n_{l_2}}=\left(2,-3,5\right)
\end{array}\right."

Then,

"\overrightarrow{n}=\overrightarrow{n_{l_1}}\times\overrightarrow{n_{l_2}}-\text{direction vector}\\[0.3cm]"



"\overrightarrow{n}=\left|\begin{array}{ccc}
\overrightarrow{i}&\overrightarrow{j}&\overrightarrow{k}\\[0.3cm]
-5&1&-2\\[0.3cm]
2&-3&5
\end{array}\right|=\\[0.3cm]
=\overrightarrow{i}\cdot\left(5-6\right)-\overrightarrow{j}\cdot\left(-25+4\right)+\overrightarrow{k}\cdot\left(15-2\right)=\\[0.3cm]
=-1\cdot\overrightarrow{i}+21\cdot\overrightarrow{j}+13\cdot\overrightarrow{k}\\[0.3cm]
\boxed{\overrightarrow{n}=\left(-1,21,13\right)}\\[0.3cm]
\left\{\begin{array}{l}
p.M\left(0,-1,-2\right)\\[0.3cm]
\overrightarrow{n}=\left(-1,21,13\right)
\end{array}\right.\longrightarrow
\left\{\begin{array}{l}
x=-t\\[0.3cm]
y=-1+21t\\[0.3cm]
z=-2+13t
\end{array}\right."

ANSWER



"\left\{\begin{array}{l}
x=-t\\[0.3cm]
y=-1+21t\\[0.3cm]
z=-2+13t
\end{array}\right."